# Token Dynamics: Towards Efficient and Dynamic Video Token Representation for Video Large Language Models

### v1

@@ Abstract
Token-based video representation has emerged as a promising approach for enabling large language models (LLMs) to interpret video content. However, existing token reduction techniques, such as token pruning and token merging, often disrupt essential spatial-temporal positional embeddings, failing to adequately balance computational efficiency with fewer tokens. Consequently, these methods result in relatively lengthy token sequences, limiting their applicability in scenarios requiring extreme token compression, such as video large language models. In this paper, we introduce the novel task of extreme short token reduction, aiming to represent extensive video sequences with a minimal number of tokens. To address this challenge, we propose Token Dynamics, a new video representation framework that dynamically reduces token count while preserving spatial-temporal coherence. Specifically, we disentangle video representations by separating visual embeddings from grid-level motion information, structuring them into: (i) a concise token base, created by clustering tokens that describe object-level content, and (ii) a token dynamics map, capturing detailed spatial-temporal motion patterns across grids. Furthermore, we introduce a cross-dynamics attention mechanism that integrates motion features into the token base without increasing token length, thereby maintaining compactness and spatial-temporal integrity. We evaluate Token Dynamics on the proposed extreme short token reduction task, demonstrating a reduction of token count to merely 0.07% of the original tokens, with only a minor performance drop of 1.13% on the NextQA-MC benchmark. Our approach similarly maintains competitive performance on ActNet-QA, Long Video Bench, and VideoMME datasets in a plug-and-play manner. Additionally, we propose two novel subtasks within extreme token reduction (fixed-length and adaptive-length compression) both effectively representing long token sequences for video-language tasks. Our method achieves state-of-the-art performance across these subtasks, offering significantly lower theoretical complexity, fewer tokens, and enhanced throughput, thus providing an efficient solution for video large language models.
@@ 1 Introduction
To address this, we introduce a structured token map representation, organized as Temporal $\times$ Width Coordinates $\times$ Height Dimension, to explicitly record spatial-temporal positions before clustering. To incorporate these dense maps without additional computational overhead, we propose a cross-dynamics attention mechanism that embeds motion information into clustered token sequences while maintaining their compact structure. Given video data’s inherent redundancy, tokens effectively represent grid-level visual information derived from pixels, offering a lightweight alternative to computationally intensive motion-tracking methods such as optical flow. Tracking token movements through the token map efficiently captures object grid-level dynamics, preserving essential spatial-temporal context with minimal computational cost.
@@
    We introduce a novel video representation framework featuring fixed/adaptive token clustering. It constructs a compact token base to capture visual pattern and object part descriptions and a token dynamics map to model their movements. A cross-motion dynamics attention mechanism integrates spatial-temporal positional information without increasing token count.
@@ 2.1 Video Large Language Models
Video Large Language Models (vLLMs) have emerged as powerful tools for bridging video understanding and natural language processing, enabling complex interpretations of video content through language-based interactions. Recent advancements have demonstrated remarkable capabilities in aligning visual and linguistic modalities, exemplified by frameworks such as LLaVA [15, 12, 14, 16], Flamingo [1], and MERLOT Reserve [29]. These methods typically rely on extensive pre-training using large-scale datasets like HD-VILA [26], InternVid [23], and NextQA [25], generating lengthy token sequences to represent videos effectively.
@@ 2.2 Token Reduction
Token reduction techniques have gained increasing attention as a means to enhance computational efficiency in Vision Transformers (ViTs) [6]. Early influential methods, such as Token Pruning [10] and Token Merging (ToMe) [3], significantly reduce computational load by identifying and removing redundant tokens or merging similar tokens, respectively. More recently, Video Token Merging (Vid-ToMe) [11] has extended ToMe to video inputs, considering temporal redundancy by merging tokens across consecutive frames. Despite these advancements, existing token reduction strategies generally adopt fixed token reduction ratios (e.g., 50%), limiting flexibility and adaptability. Such fixed strategies can either inadequately reduce redundant tokens, resulting in lingering inefficiencies, or inadvertently merge tokens representing distinct objects, thus losing crucial spatial-temporal dynamics necessary for precise video interpretation. To overcome these limitations, our proposed Token Dynamics framework introduces dynamic token clustering to generate a compact token representation while explicitly preserving spatial-temporal motion information via a dedicated token dynamics map. Through our novel cross-dynamics attention mechanism, our approach effectively integrates spatial-temporal coherence into concise token sequences without compromising accuracy, outperforming state-of-the-art token reduction methods, even in scenarios of extreme token compression.
Figure 1: Overview of the Token Dynamics framework and its integration into video large language models. The input video is first tokenized into a 1D token sequence and subsequently undergoes token clustering to group similar tokens, potentially representing the same object parts at the grid level. This clustering yields our proposed Token Dynamics representation, comprising two complementary components: (i) a concise token base, dynamically sized by merging tokens into representative cluster centroids, and (ii) a token dynamics map, a three-dimensional tensor explicitly capturing grid-level spatial-temporal coordinates (i.e., x, y, and frame index as spatial-temporal positions across frames) to represent detailed motion dynamics for each cluster. To efficiently integrate motion information into the token base, the token dynamics map is first flattened and dimensionally aligned via an MLP, and then processed by a cross-dynamics attention module employing two multi-head attention operations. Crucially, this integration preserves the compactness of the original token count. Finally, the compressed token sequence, along with a tokenized query, is projected into the large language model for zero-shot inference.
@@ 3.3 Token Dynamics Representation
We introduce the Token Dynamics Representation, designed to minimize token sequence length dynamically while preserving complete spatial-temporal motion information. We conceptualize videos as compositions of multiple scenes, each containing objects with distinct attributes and motions. To achieve this, we reformulate the original video representation (token sequences from Vision Transformers [6]) into two components: (i) a concise token base, representing objects in each scene as token centroids obtained by clustering similar visual patterns at the grid-level; (ii) a token dynamics map, explicitly recording the spatial-temporal coordinates of each cluster centroid, capturing detailed object-level grid trajectories. As illustrated in Fig. 1, this design ensures the token base and dynamics map are disentangled at the grid-level. Thus, we represent video content concisely without sacrificing crucial spatial-temporal information, achieving efficiency in token sequence reduction.
@@ 3.3.1 Token Clustering
3.3.2 Concise Token Base
To construct a compact representation while preserving essential visual patterns and object part information, we merge clustered tokens into representative centroids, forming the concise token base. Using the token indices from Equation 2, the $k$-th token base $b_{k}$ is computed as:
@@ 3.3.2 Concise Token Base
3.3.3 Token Dynamics Map
While the concise token base effectively represents visual patterns, it lacks spatial-temporal details essential for video understanding. To preserve motion dynamics, we record spatial-temporal coordinates during clustering, constructing a detailed motion representation—i.e., the token dynamics map.
Specifically, we formulate a three-dimensional tensor capturing $x$, $y$, and frame indices, ensuring that each cluster centroid retains its original spatial-temporal position. This map aligns with the downsampled video structure from the ViT and is defined as:
@@ 3.4 Cross Dynamics Attention Module
While the token base and token dynamics map effectively disentangle visual and motion information, directly embedding the token dynamics map would significantly increase computational overhead due to its large size. To address this, we propose a cross-dynamics attention module that integrates motion dynamics into the concise token base without increasing token count or dimensionality.
First, the token dynamics map is projected via an MLP to align its dimensions with the token base. Then, a multi-head transformer decoder extracts motion information by treating the projected token dynamics map as encoded features and the token base as queries. This process embeds spatial-temporal dynamics into the token sequence while maintaining its compact structure:
@@
where $\mathcal{F}_{\text{A}}$ represents multi-head attention, $m$ is the token dynamics map, $b$ is the concise token base, and $W_{r}$ denotes the weight matrix in the $r$-th MLP layer. Superscripts indicate tensor shapes. The output $b_{\text{bank}}$ is the refined compressed token sequence, which is subsequently fed into the LLM for downstream tasks.
@@ 4.5 Performance on Multiple Benchmarks
Table 6: Ablation study on the impact of individual submodules on model performance. We incrementally add each proposed module component to the original baseline, following Sec. 3. A ✓indicates that the corresponding submodule is applied, while R denotes that its parameters are randomized. Note: For the ablation of w/. Token base + map, cross-dynamics attention parameters are randomized, as the vLLM cannot directly utilize the token dynamics map.
@@ 4.6 Ablation Study
Quantitative Ablation. To assess the contribution of each module, we conduct an ablation study by incrementally adding components and evaluating their impact. In Table 6, Base, Map, and Attn denote the token base (Sec. 3.3.2), token dynamics map (Sec. 3.3.3), and cross-dynamics attention (Sec. 3.4), respectively. Using only the Base reduces tokens to 0.14% of the original but results in a severe 22.0% accuracy drop, emphasizing the importance of spatial-temporal information. Adding the Map slightly improves accuracy but remains ineffective as the LLM cannot directly utilize motion cues. When cross-dynamics attention is introduced, spatial-temporal information is successfully integrated, achieving extreme token reduction while significantly restoring accuracy. To further verify module necessity, we randomize their parameters (R), leading to substantial accuracy drops, reinforcing their critical role in maintaining representation quality. This study highlights that while token clustering compresses redundant information, preserving motion dynamics via the token map and cross-dynamics attention is essential for maintaining accuracy in extreme token reduction settings.
Visualization of Token Clusters. As shown in Fig. 3, each object can be associated with one or multiple clusters and may be represented by either a single token or multiple tokens, as visual tokenizers act as descriptors based on similarity in token space. To analyze token clustering effectiveness, we compare an adaptive K-Means clustering approach with the state-of-the-art segmentation model Segment Anything (SAM) [20]. Specifically, adaptive K-Means is directly applied to token sequences, while the segmentation method assigns cluster indices based on the number of detected segmentation masks. While SAM achieves more refined clustering, both approaches successfully group similar objects and maintain cluster consistency across frames. Given the computational efficiency requirements of video large language models (vLLMs), adaptive K-Means provides a sufficiently precise clustering mechanism, effectively tracking object motions and trajectories for the subsequent token dynamics map.
@@ 5 Conclusion
In this work, we introduced Token Dynamics, a novel video representation framework designed to achieve extreme token reduction while preserving essential spatial-temporal coherence for video large language models (vLLMs). Our approach reformulates video representation by disentangling visual embeddings and motion information into two complementary components: (i) a concise token base that clusters object-descriptive tokens, and (ii) a token dynamics map that captures fine-grained spatial-temporal motion patterns. We further proposed a cross-dynamics attention mechanism that integrates motion features into the token base without increasing token length, ensuring efficient and compact video representations. To systematically evaluate our framework, we introduced the novel Extreme Short Token Reduction task, encompassing both fixed-length and adaptive-length token compression settings. Our method demonstrated strong performance across multiple video-language benchmarks, reducing token count to as little as 0.07% of the original sequence while maintaining competitive accuracy. Specifically, we achieved only a minor 1.13% accuracy drop on the NextQA-MC benchmark and maintained comparable performance on ActNet-QA, Long Video Bench, and VideoMME, validating the effectiveness of our approach. Extensive ablation studies further highlighted the contributions of each module, confirming that both token base clustering and motion integration via cross-dynamics attention are critical for preserving video understanding. Moreover, efficiency analyses demonstrated the practical benefits of Token Dynamics, significantly reducing computational overhead compared to existing token reduction methods. Our framework is designed as a plug-and-play solution, making it broadly applicable to existing vLLMs without requiring architectural modifications. By balancing extreme token compression with high accuracy retention, Token Dynamics provides a scalable and efficient solution for video-language models, particularly in resource-constrained environments. Future work will explore further optimizations for large-scale models, extend token dynamics to more complex tasks like video generation and trajectory prediction.

### v2

@@ Abstract
Token-based video representation has emerged as a promising approach for enabling large language models (LLMs) to interpret video content. However, existing token reduction techniques, such as token pruning and token merging, often disrupt essential spatial-temporal positional embeddings, failing to adequately balance computational efficiency with fewer tokens. Consequently, these methods result in relatively lengthy token sequences, limiting their applicability in scenarios requiring extreme token compression, such as video large language models. In this paper, we introduce the novel task of extreme short token reduction, aiming to represent extensive video sequences with a minimal number of tokens. To address this challenge, we propose Token Dynamics, a new video representation framework that dynamically reduces token count while preserving spatial-temporal coherence. Specifically, we disentangle video representations by separating visual embeddings from grid-level motion information, structuring them into: (i) a concise token hash table, created by clustering tokens that describe object-level content, (ii) a token indices key map, capturing detailed spatial-temporal motion patterns across grids, (iii) a token hash function, which vector-quantizes the token hash table to reconstruct the token sequence from the key map. Furthermore, we introduce a cross-dynamics attention mechanism that integrates motion features into the token hash table without increasing token length, thereby maintaining compactness and spatial-temporal integrity. We evaluate Token Dynamics on the proposed extreme short token reduction task, demonstrating a reduction of token count to merely 0.07% of the original tokens, with only a minor performance drop of 1.13% on the NextQA-MC benchmark. Our approach similarly maintains competitive performance on ActNet-QA, Long Video Bench, and VideoMME datasets in a plug-and-play manner. Additionally, we propose two novel subtasks within extreme token reduction (fixed-length and adaptive-length compression) both effectively representing long token sequences for video-language tasks. Our method achieves state-of-the-art performance across these subtasks, offering significantly lower theoretical complexity, fewer tokens, and enhanced throughput, thus providing an efficient solution for video large language models.
@@ 1 Introduction
To address this, we introduce a structured token indices key map representation, organized as Temporal $\times$ Width Coordinates $\times$ Height Dimension, to explicitly record spatial-temporal positions before clustering. To incorporate these dense maps without additional computational overhead, we propose a cross-dynamics attention mechanism that embeds motion information into clustered token sequences while maintaining their compact structure. Given video data’s inherent redundancy, tokens effectively represent grid-level visual information derived from pixels, offering a lightweight alternative to computationally intensive motion-tracking methods such as optical flow. Tracking token movements through the token indices key map efficiently captures object grid-level dynamics, preserving essential spatial-temporal context with minimal computational cost.
@@
    We introduce a novel video representation framework featuring fixed/adaptive token clustering. It constructs a compact token hash table vector to capture visual pattern and object part descriptions and a token indices key map to model their movements. A cross-motion dynamics attention mechanism integrates spatial-temporal positional information without increasing token count.
@@ 2.1 Video Large Language Models
Video Large Language Models (vLLMs) have emerged as powerful tools for bridging video understanding and natural language processing, enabling complex interpretations of video content through language-based interactions. Recent advancements have demonstrated remarkable capabilities in aligning visual and linguistic modalities, exemplified by frameworks such as LLaVA [16, 13, 15, 17], Flamingo [1], AuroraCap [5], and MERLOT Reserve [30]. These methods typically rely on extensive pre-training using large-scale datasets like HD-VILA [27], InternVid [24], and NextQA [26], generating lengthy token sequences to represent videos effectively.
@@ 2.2 Token Reduction
Token reduction techniques have gained increasing attention as a means to enhance computational efficiency in Vision Transformers (ViTs) [7]. Early influential methods, such as Token Pruning [11] and Token Merging (ToMe) [3], significantly reduce computational load by identifying and removing redundant tokens or merging similar tokens, respectively. More recently, Video Token Merging (Vid-ToMe) [12] has extended ToMe to video inputs, considering temporal redundancy by merging tokens across consecutive frames. Despite these advancements, existing token reduction strategies generally adopt fixed token reduction ratios (e.g., 50%), limiting flexibility and adaptability. Such fixed strategies can either inadequately reduce redundant tokens, resulting in lingering inefficiencies, or inadvertently merge tokens representing distinct objects, thus losing crucial spatial-temporal dynamics necessary for precise video interpretation. To overcome these limitations, our proposed Token Dynamics framework introduces dynamic token clustering to generate a compact token representation while explicitly preserving spatial-temporal motion information via a dedicated token indices key map. Through our novel cross-dynamics attention mechanism, our approach effectively integrates spatial-temporal coherence into concise token sequences without compromising accuracy, outperforming state-of-the-art token reduction methods, even in scenarios of extreme token compression.
Figure 1: Overview of the Token Dynamics framework and its integration into video large language models. The input video is first tokenized into a 1D token sequence and subsequently undergoes token clustering to group similar tokens, potentially representing the same object parts at the grid level. This clustering yields our proposed Token Dynamics representation, comprising two complementary components: (i) a concise token hash table, dynamically sized by merging tokens into representative cluster centroids, and (ii) a token indices key map, a three-dimensional tensor explicitly capturing grid-level spatial-temporal coordinates (i.e., x, y, and frame index as spatial-temporal positions across frames) to represent detailed motion dynamics for each cluster. To efficiently integrate motion information into the token hash table, the token indices key map is first flattened and dimensionally aligned via an MLP, and then processed by a cross-dynamics attention module employing two multi-head attention operations. Crucially, this integration preserves the compactness of the original token count. Finally, the compressed token sequence, along with a tokenized query, is projected into the large language model for zero-shot inference.
@@ 3.3 Token Dynamics Representation
We introduce the Token Dynamics Representation, designed to minimize token sequence length dynamically while preserving complete spatial-temporal motion information. We conceptualize videos as compositions of multiple scenes, each containing objects with distinct attributes and motions. To achieve this, we reformulate the original video representation (token sequences from Vision Transformers [7]) into two components: (i) a concise token hash table, representing objects in each scene as token centroids obtained by clustering similar visual patterns at the grid-level; (ii) a token indices key map, explicitly recording the spatial-temporal coordinates of each cluster centroid, capturing detailed object-level grid trajectories. As illustrated in Fig. 1, this design ensures the token hash table and dynamics map are disentangled at the grid-level. Thus, we represent video content concisely without sacrificing crucial spatial-temporal information, achieving efficiency in token sequence reduction.
@@ 3.3.1 Token Clustering
3.3.2 Concise Token Hash Table
To construct a compact representation while preserving essential visual patterns and object part information, we merge clustered tokens into representative centroids, forming the concise token hash table. Using the token indices from Equation 2, the $k$-th token hash table $b_{k}$ vector is computed as:
@@ 3.3.2 Concise Token Hash Table
3.3.3 Token Indices Key Map
While the concise token hash table effectively represents visual patterns, it lacks spatial-temporal details essential for video understanding. To preserve motion dynamics, we record spatial-temporal coordinates during clustering, constructing a detailed motion representation—i.e., the token indices key map.
Specifically, we formulate a three-dimensional tensor capturing $x$, $y$, and frame indices, ensuring that each cluster centroid retains its original spatial-temporal position. This map aligns with the downsampled video structure from the ViT and is defined by a token hash function:
@@ 3.4 Cross Dynamics Attention Module
While the token hash table and token indices key map effectively disentangle visual and motion information, directly embedding the token indices key map would significantly increase computational overhead due to its large size. To address this, we propose a cross-dynamics attention module that integrates motion dynamics into the concise token hash table without increasing token count or dimensionality.
First, the token indices key map is projected via an MLP to align its dimensions with the token hash table. Then, a multi-head transformer decoder extracts motion information by treating the projected token indices key map as encoded features and the token hash table as queries. This process embeds spatial-temporal dynamics into the token sequence while maintaining its compact structure:
@@
where $\mathcal{F}_{\text{A}}$ represents multi-head attention, $m$ is the token indices key map, $b$ is the concise token hash table, and $W_{r}$ denotes the weight matrix in the $r$-th MLP layer. Superscripts indicate tensor shapes. The output $b_{\text{bank}}$ is the refined compressed token sequence, which is subsequently fed into the LLM for downstream tasks.
@@ 4.5 Performance on Multiple Benchmarks
Table 6: Ablation study on the impact of individual submodules on model performance. We incrementally add each proposed module component to the original baseline, following Sec. 3. A ✓indicates that the corresponding submodule is applied, while R denotes that its parameters are randomized. Note: For the ablation of w/. Token Hash Table + map, cross-dynamics attention parameters are randomized, as the vLLM cannot directly utilize the token indices key map.
@@ 4.6 Ablation Study
Quantitative Ablation. To assess the contribution of each module, we conduct an ablation study by incrementally adding components and evaluating their impact. In Table 6, Base, Map, and Attn denote the token hash table (Sec. 3.3.2), token indices key map (Sec. 3.3.3), and cross-dynamics attention (Sec. 3.4), respectively. Using only the Base reduces tokens to 0.14% of the original but results in a severe 22.0% accuracy drop, emphasizing the importance of spatial-temporal information. Adding the Map slightly improves accuracy but remains ineffective as the LLM cannot directly utilize motion cues. When cross-dynamics attention is introduced, spatial-temporal information is successfully integrated, achieving extreme token reduction while significantly restoring accuracy. To further verify module necessity, we randomize their parameters (R), leading to substantial accuracy drops, reinforcing their critical role in maintaining representation quality. This study highlights that while token clustering compresses redundant information, preserving motion dynamics via the token map and cross-dynamics attention is essential for maintaining accuracy in extreme token reduction settings.
Visualization of Token Clusters. As shown in Fig. 3, each object can be associated with one or multiple clusters and may be represented by either a single token or multiple tokens, as visual tokenizers act as descriptors based on similarity in token space. To analyze token clustering effectiveness, we compare an adaptive K-Means clustering approach with the state-of-the-art segmentation model Segment Anything (SAM) [21]. Specifically, adaptive K-Means is directly applied to token sequences, while the segmentation method assigns cluster indices based on the number of detected segmentation masks. While SAM achieves more refined clustering, both approaches successfully group similar objects and maintain cluster consistency across frames. Given the computational efficiency requirements of video large language models (vLLMs), adaptive K-Means provides a sufficiently precise clustering mechanism, effectively tracking object motions and trajectories for the subsequent token indices key map.
@@ 5 Conclusion
In this work, we introduced Token Dynamics, a novel video representation framework designed to achieve extreme token reduction while preserving essential spatial-temporal coherence for video large language models (vLLMs). Our approach reformulates video representation by disentangling visual embeddings and motion information into two complementary components: (i) a concise token hash table that clusters object-descriptive tokens, and (ii) a token indices key map that captures fine-grained spatial-temporal motion patterns. We further proposed a cross-dynamics attention mechanism that integrates motion features into the token hash table without increasing token length, ensuring efficient and compact video representations. To systematically evaluate our framework, we introduced the novel Extreme Short Token Reduction task, encompassing both fixed-length and adaptive-length token compression settings. Our method demonstrated strong performance across multiple video-language benchmarks, reducing token count to as little as 0.07% of the original sequence while maintaining competitive accuracy. Specifically, we achieved only a minor 1.13% accuracy drop on the NextQA-MC benchmark and maintained comparable performance on ActNet-QA, Long Video Bench, and VideoMME, validating the effectiveness of our approach. Extensive ablation studies further highlighted the contributions of each module, confirming that both token hash table clustering and motion integration via cross-dynamics attention are critical for preserving video understanding. Moreover, efficiency analyses demonstrated the practical benefits of Token Dynamics, significantly reducing computational overhead compared to existing token reduction methods. Our framework is designed as a plug-and-play solution, making it broadly applicable to existing vLLMs without requiring architectural modifications. By balancing extreme token compression with high accuracy retention, Token Dynamics provides a scalable and efficient solution for video-language models, particularly in resource-constrained environments. Future work will explore further optimizations for large-scale models, extend token dynamics to more complex tasks like video generation and trajectory prediction.